# Scale-Dependent Multifractality in Bitcoin Realised Volatility: Implications for Rough Volatility Modelling

### v1

@@ 3.2 Data Description and Preprocessing
To ensure consistency and minimise missing data effects, for each calendar year $Y$, we systematically selected the most complete 90-day window of consecutive minute-level BTC/USD observations. We opted to analyse one representative 90-day high-quality segment per year to control for seasonality and to ensure stationarity over each segment. Using longer or multiple windows risked including structural changes or requiring imputation for extensive gaps. Let $T=129{,}600$ minutes denote the target window length. The search space was defined as the set of all candidate windows of length $T$, starting at time $t_{s}$ and ending at $t_{e}=t_{s}+T-1$, with $t_{s}$ incremented in daily steps (1,440 minutes). For each candidate window, we constructed the corresponding sequence of timestamps $\{t_{i}\}_{i=1}^{T}$ and computed the number of missing Close price observations $m(t_{s})$. The optimal window $[t_{s}^{*},t_{e}^{*}]$ was selected as:
@@ 3.5 Final Estimation Setup
[…]
@@ 4.1 Main Result: Roughness Estimation behaviour
As shown in Table 3, $\log W(L,K,p)$ remains strictly negative across all years, frequencies, and both the standard and extended $p$-grids. No zero-crossing is observed in any configuration, with maximum values not exceeding $-0.009$, even under the wide grid. This behaviour deviates from the expected pattern under homogeneous roughness, where a unique zero-crossing at $p=\hat{p}$ would yield a well-defined roughness exponent $H=1/\hat{p}$. The consistent absence of such a crossing indicates a systematic failure of the normalised $p$-variation estimator in the Bitcoin setting. The robustness of this result across years and frequencies suggests it reflects a structural property rather than a data artefact.
@@ 6 Conclusion
Table A.2: Summary of missing data and large gaps ($>$1 hour) by frequency and year in retained 90-day BTC windows.
Frequency Year Total Missing Minutes Gaps $>$1 Hour (Count)
@@
Frequency Year $N_{\text{obs}}$Breaks Break Indices

### v2

@@ 3.2 Data Description and Preprocessing
To ensure consistency and minimise missing data effects, for each calendar year $Y$, we systematically selected the most complete 90-day window of consecutive minute-level BTC/USD observations. We opted to analyse one representative 90-day high-quality segment per year to control for seasonality and to ensure stationarity over each segment. Using longer or multiple windows risked including structural changes or requiring imputation for extensive gaps. Let T=129,600T=129{,}600 minutes denote the target window length. The search space was defined as the set of all candidate windows of length $T$, starting at time $t_{s}$ and ending at $t_{e}=t_{s}+T-1$, with $t_{s}$ incremented in daily steps (1,440 minutes). For each candidate window, we constructed the corresponding sequence of timestamps $\{t_{i}\}_{i=1}^{T}$ and computed the number of missing Close price observations $m(t_{s})$. The optimal window $t_{s}^{*},t_{e}^{*}]$ was selected as:
@@ 3.3 Realised Volatility Construction
A potential concern, highlighted in Fukasawa et al., (2022), is that realized volatility constructed from single low-frequency returns
$RV^{\Delta}_{t}=|r_{t,\Delta}|$
may be contaminated by white-noise components, leading to artificially rough behaviour. To address this, we also consider a noise-robust construction: for each frequency $\Delta$, one realized volatility observation is computed as
$\widetilde{RV}^{\Delta}_{t}=\left(\sum_{i=1}^{\Delta}r_{t,i}^{2}\right)^{1/2},$
where $\{r_{t,i}\}$ are one-minute log-returns within the $\Delta$-minute interval. This definition ensures that each volatility estimate incorporates multiple high-frequency increments, thereby mitigating the influence of white-noise domination. As reported in Appendix A.12, the use of $\widetilde{RV}^{\Delta}_{t}$ does not alter our main findings: the statistic $W(p)$ remains strictly negative across all years and frequencies, indicating that the absence of a zero-crossing cannot be explained by noise contamination alone.
@@ 3.5 Final Estimation Setup
For each $(\text{year},\text{frequency})$ pair, we use the optimal block parameters $(K,n)$ determined in Section 3.4. The block sums $S_{j}$ are computed, and the normalised $p$-variation statistic is evaluated as:
@@ 4.1 Main Result: Roughness Estimation behaviour
As shown in Table 3, $\log W(L,K,p)$ remains strictly negative across all years, frequencies, and both the standard and extended $p$-grids. No zero-crossing is observed in any configuration, with maximum values not exceeding 0.009-0.009, even under the wide grid. This behaviour deviates from the expected pattern under homogeneous roughness, where a unique zero-crossing at $p=\hat{p}$ would yield a well-defined roughness exponent $H=1/\hat{p}$. The consistent absence of such a crossing indicates a systematic failure of the normalised $p$-variation estimator in the Bitcoin setting. The robustness of this result across years and frequencies suggests it reflects a structural property rather than a data artefact.
@@ 6 Conclusion
Table A.2: Summary of missing data and large gaps (>>1 hour) by frequency and year in retained 90-day BTC windows.
Frequency Year Total Missing Minutes Gaps >>1 Hour (Count)
@@
Frequency Year $N_{\text{obs}}$ #Breaks Break Indices
@@
Table A.12: Minimum and maximum values of $\log W(L,K,p)$ for each (frequency, year) pair, using the noise-robust realised volatility $\widetilde{RV}^{\Delta}_{t}$.
Frequency Year Standard Grid Wide Grid
Min Max Min Max
1min 2017 -1.797 -0.090 -1.797 -0.009
1min 2018 -3.454 -0.110 -3.454 -0.011
1min 2019 -3.149 -0.114 -3.149 -0.011
1min 2020 -1.857 -0.124 -1.857 -0.012
1min 2021 -2.377 -0.098 -2.377 -0.010
1min 2022 -5.021 -0.154 -5.021 -0.015
1min 2023 -6.177 -0.199 -6.177 -0.020
1min 2024 -5.249 -0.175 -5.249 -0.018
5min 2017 -1.188 -0.067 -1.188 -0.007
5min 2018 -2.738 -0.089 -2.738 -0.009
5min 2019 -2.625 -0.093 -2.625 -0.009
5min 2020 -1.434 -0.104 -1.434 -0.010
5min 2021 -1.855 -0.079 -1.855 -0.008
5min 2022 -4.442 -0.133 -4.442 -0.013
5min 2023 -5.658 -0.176 -5.658 -0.018
5min 2024 -4.778 -0.155 -4.778 -0.016
10min 2017 -1.093 -0.062 -1.093 -0.006
10min 2018 -2.628 -0.084 -2.628 -0.008
10min 2019 -2.542 -0.089 -2.542 -0.009
10min 2020 -1.388 -0.100 -1.388 -0.010
10min 2021 -1.782 -0.075 -1.782 -0.008
10min 2024 -4.831 -0.150 -4.831 -0.015
15min 2018 -2.588 -0.083 -2.588 -0.008
15min 2019 -2.446 -0.087 -2.446 -0.009
15min 2020 -1.402 -0.099 -1.402 -0.010
15min 2021 -1.908 -0.073 -1.908 -0.007
15min 2024 -4.807 -0.150 -4.807 -0.015